# itemizestditemize

### v1

@@ Abstract
Figure 1: CRISP computes entirely new contact sequences from naive and even all-zero initializations. For (a), (b), (c), and (d), the left side shows the initial trajectories, and the right side displays the optimized trajectory from CRISP. For (e) the hopper problem, the initial guess is a free-fall motion released from the origin. The color gradient represents the progression of time (from blue to yellow). Detailed descriptions of these tasks are provided in §IV.
@@ I Introduction
Contact-implicit motion planning. Among the many efforts for motion planning through contact (see §V for a review), the so-called contact-implicit formulation stood out as a particularly popular and promising approach [45]. This formulation distinguishes itself by integrating contact dynamics into the motion planning framework through the introduction of complementarity constraints, which allow simultaneous optimization of the state-control trajectories and the contact forces without explicit (and discrete) mode switching (cf. (8)). However, the contact-implicit formulation circumvents explicit mode switching at the price of arriving at an optimization problem known as Mathematical Programming with Complementarity Constraints (MPCC). Although they “look like” smooth and continuous optimization problems, MPCC problems are known to fail the classical constraint qualifications (CQs) at all feasible points (e.g., Linear Independence Constraint Qualification and Mangasarian-Fromovitz Constraint Qualification) [46, 47, 58]. To make this concrete, we provide the geometric intuition through a simple MPCC problem.
@@ IV-B Benchmark Problems
    Hopper: We model the system as a point mass $m$ atop a massless leg with original length $l_{0}$, capable of radial contraction and thrust application. As shown in Fig. 4(d), the hopper exhibits two distinct phases: the “Flight Phase”, where the leg length remains constant and motion approximates free fall, and the “Stance Phase”, which initiates upon ground contact. During flight, we directly control the leg angular speed ($\dot{\theta}=u_{1}$). The stance phase, triggered by $p_{y}-l_{0}\cos\theta\leq 0$, involves fixed-point rotation around the contact point. As the leg contracts ($r$ increases within $r_{0}$), it can exert a unidirectional thrust $u=u_{2}$ along its radial axis. The transition back to flight occurs when $r=0$. The task objective is to navigate the hopper’s hybrid dynamics without pre-specified contact sequences. The hopper is released from a certain height at the origin, and must optimize its own contact sequence, stance phase contact forces, and flight phase angular velocities to jump and stop at $x=2\,\textup{m}$. This involves simultaneously managing multiple goals: exploring optimal contact sequences, controlling leg angle during flight, and managing thrust after landing to achieve the specified final position. The initial guess is a passive free-fall motion. As derived in Appendix B-D, we employ a contact-implicit formulation to unify the dynamics of both phases within a single framework, with carefully crafted complementarity constraints to govern the transitions between phases.
@@
    Waiter: The waiter problem, see Fig. 4(e), is also studied in [57] where it was modeled with linear complementarity constraints and optimized in a model-predictive control pipeline. Initially, the plate on the table has only a small overhang. The objective is to indirectly control the friction force with the plate by applying a normal force $\lambda$ and a pulling force $u$ to the plate $m_{1}$ via the pusher $m_{2}$. The goal is to gradually extract the plate until its center of mass $x_{1}$ aligns with the pusher’s center $x_{2}$, stopping at the table’s edge with matching velocities. This task requires consideration of the friction between the table and the plate, as well as preventing the plate from tipping over its leftmost contact point with the table (see Appendix B-E). We provide an all-zero initial guess to test the solvers’ ability to optimize an informed contact trajectory.
@@ IV-C Numerical Results
Methods Cartpole with Soft Walls Push Box Transport
Success Rate [%] Tracking Error Violation Iterations Time [s] Success Rate [%] Tracking Error Violation Iterations Time [s] Success Rate [%] Tracking Error Violation Iterations Time [s]
Snopt $90$ $0.08$ $2.0\times 10^{-3}$ $4552.50$ $12.97$ $100$ $4.25$ $1.0\times 10^{-11}$ $3303.00$ $25.22$ $\bf{100}$ $1.7\times 10^{-4}$ $5.5\times 10^{-9}$ $5782.30$ $43.98$
Ipopt $96.67$ $0.04$ $8.9\times 10^{-7}$ $\bf 288.21$ $\bf 0.51$ $\bf{100}$ $0.03$ $\bf 1.3\times 10^{-11}$ $\bf 407.69$ $0.92$ $\bf{100}$ $0.02$ $1.6\times 10^{-8}$ $404.96$ $1.36$
PROXNLP $70$ $0.032$ $1.4\times 10^{-3}$ $716.25$ $140.92$ $66.67$ $0.03$ $3.6\times 10^{-5}$ $2000.00$ $1277.21$ $59.26$ $0.01$ $1.1\times 10^{-3}$ $361.92$ $1070.01$
CRISP (ours) $\bf{100}$ $\bf 0.02$ $\bf 1.4\times 10^{-7}$ $415.39$ $1.56$ $\bf{100}$ $\bf 0.02$ $8.3\times 10^{-10}$ $420.69$ $\bf{0.74}$ $\bf{100}$ $\bf 1.1\times 10^{-4}$ $\bf 1.1\times 10^{-11}$ $\bf{286.20}$ $\bf 0.77$
@@
Figure 5: Box plots of benchmark metrics on three tasks: cartpole with soft walls, push box, and transport. The plots complement  Table II to show the distribution of (a) objective value, (b) tracking error, and (c) constraint violation across multiple initial states and initial guesses.
@@
Quantitative results. For the first three examples, we recorded and summarized the numerical performance of the benchmark solvers under different initial states and initial guesses, as presented in the Table II and Fig. 5. We set a constraint violation threshold of $1\times 10^{-6}$ for each solver when running the experiments, and all final results with feasibility below $1\times 10^{-5}$ were considered successful, with their average constraint violation recorded. The objective function, as previously highlighted, primarily balances tracking error and control efforts. We analyze the performance of four solvers across three examples, focusing on median tracking error, average iteration count, and computation time. To better illustrate the performance and robustness of each solver under different initial guesses, we boxplot the distribution of constraint violations, objective values, and tracking errors in Fig. 5. These plots provide a comprehensive overview of solver behavior across all experiments, including both feasible and infeasible solutions, offering a complete picture of their performance characteristics.
@@ B-C Transport
B-D Hopper
@@ Appendix C Visualization of Numerical Results
Figure 8: Visualization of more trajectories solved by CRISP in cartpole with soft walls (row 1-2), push box (row 3-4), and payload transport (row 5-6) under different initial conditions.

### v2

@@ Abstract
Figure 1: CRISP computes entirely new contact sequences from naive and even all-zero initializations. For (a), (b), (c), and (d), the left side shows the initial trajectories and the right side displays the optimized trajectory from CRISP. For (e) the hopper problem, the initial guess is a free-fall motion released from the origin. The color gradient represents the progression of time (from blue to yellow). For (f), we implement the policy derived from CRISP in a Model Predictive Control (MPC) framework for real-world push tasks. Detailed descriptions of these tasks are provided in §IV.
@@ I Introduction
Contact-implicit motion planning. Among the many efforts for motion planning through contact (see §V for a review), the so-called contact-implicit formulation stood out as a particularly popular and promising approach [45, 59]. This formulation distinguishes itself by integrating contact dynamics into the motion planning framework through the introduction of complementarity constraints, which allow simultaneous optimization of the state-control trajectories and the contact forces without explicit (and discrete) mode switching (cf. (8)). However, the contact-implicit formulation circumvents explicit mode switching at the price of arriving at an optimization problem known as Mathematical Programming with Complementarity Constraints (MPCC). Although they “look like” smooth and continuous optimization problems, MPCC problems are known to fail the classical constraint qualifications (CQs) at all feasible points (e.g., Linear Independence Constraint Qualification and Mangasarian-Fromovitz Constraint Qualification) [46, 47, 58]. To make this concrete, we provide the geometric intuition through a simple MPCC problem.
@@ IV-B Benchmark Problems
    Push T: The modeling techniques for the push T problem are similar to those used in the push box problem. However, it presents its own challenges due to the non-convex shapes of the “T” and the increase in contact modes ($\lambda_{1}-\lambda_{8}$ as shown in Fig. 4(d)). This means we cannot simply construct complementarity constraints related to contact using four hyperplanes as we do with the push box problem. We now need to establish complementarity conditions that involve contact points lying on certain line segments and exerting corresponding contact forces, which requires special modeling techniques. Details of the problem formulation is demonstrated in Appendix B-D. Together with the slack variables, the total state dimension is 29.
@@
    Hopper: We model the system as a point mass $m$ atop a massless leg with original length $l_{0}$, capable of radial contraction and thrust application. As shown in Fig. 4(e), the hopper exhibits two distinct phases: the “Flight Phase”, where the leg length remains constant and motion approximates free fall, and the “Stance Phase”, which initiates upon ground contact. During flight, we directly control the leg angular speed ($\dot{\theta}=u_{1}$). The stance phase, triggered by $p_{y}-l_{0}\cos\theta\leq 0$, involves fixed-point rotation around the contact point. As the leg contracts ($r$ increases within $r_{0}$), it can exert a unidirectional thrust $u=u_{2}$ along its radial axis. The transition back to flight occurs when $r=0$. The task objective is to navigate the hopper’s hybrid dynamics without pre-specified contact sequences. The hopper is released from a certain height at the origin, and must optimize its own contact sequence, stance phase contact forces, and flight phase angular velocities to jump and stop at $x=2\,\textup{m}$. This involves simultaneously managing multiple goals: exploring optimal contact sequences, controlling leg angle during flight, and managing thrust after landing to achieve the specified final position. The initial guess is a passive free-fall motion. As derived in Appendix B-E, we employ a contact-implicit formulation to unify the dynamics of both phases within a single framework, with carefully crafted complementarity constraints to govern the transitions between phases.
  6. 6.
    Waiter: The waiter problem, see Fig. 4(f), is also studied in [57] where it was modeled with linear complementarity constraints and optimized in a model-predictive control pipeline. Initially, the plate on the table has only a small overhang. The objective is to indirectly control the friction force with the plate by applying a normal force $\lambda$ and a pulling force $u$ to the plate $m_{1}$ via the pusher $m_{2}$. The goal is to gradually extract the plate until its center of mass $x_{1}$ aligns with the pusher’s center $x_{2}$, stopping at the table’s edge with matching velocities. This task requires consideration of the friction between the table and the plate, as well as preventing the plate from tipping over its leftmost contact point with the table (see Appendix B-F). We provide an all-zero initial guess to test the solvers’ ability to optimize an informed contact trajectory.
@@ IV-C Numerical Results
Methods Cartpole with Soft Walls Push Box Transport Push T
[…]
Snopt $90$ $0.08$ $2.0\times 10^{-3}$ $4552.50$ $12.97$ $\bf 100$ $4.25$ $1.0\times 10^{-11}$ $3303.00$ $25.22$ $\bf{100}$ $1.7\times 10^{-4}$ $5.5\times 10^{-9}$ $5782.30$ $43.98$ $\bf{100}$ $0.67$ $\bf 6.4\times 10^{-8}$ $2257.50$ $210.92$
Ipopt $96.67$ $0.04$ $8.9\times 10^{-7}$ $\bf 288.21$ $\bf 0.51$ $\bf{100}$ $0.03$ $\bf 1.3\times 10^{-11}$ $\bf 407.69$ $0.92$ $\bf{100}$ $0.02$ $1.6\times 10^{-8}$ $404.96$ $1.36$ $\bf{100}$ $0.26$ $4.9\times 10^{-7}$ $1560.20$ $39.49$
PROXNLP $70$ $0.032$ $1.4\times 10^{-3}$ $716.25$ $140.92$ $66.67$ $0.03$ $3.6\times 10^{-5}$ $2000.00$ $1277.21$ $59.26$ $0.01$ $1.1\times 10^{-3}$ $361.92$ $1070.01$- -- -- -- --
CRISP (ours) $\bf{100}$ $\bf 0.02$ $\bf 1.4\times 10^{-7}$ $415.39$ $1.56$ $\bf{100}$ $\bf 0.02$ $8.3\times 10^{-10}$ $420.69$ $\bf{0.74}$ $\bf{100}$ $\bf 1.1\times 10^{-4}$ $\bf 1.1\times 10^{-11}$ $\bf{286.20}$ $\bf 0.77$ $\bf{100}$ $\bf 0.01$ $8.7\times 10^{-7}$ $\bf 311.35$ $\bf 3.54$
@@
Figure 5: Box plots of benchmark metrics on three tasks: cartpole with soft walls, push box, transport, and push T. The plots complement  Table II to show the distribution of (a) objective value, (b) tracking error, and (c) constraint violation across multiple initial states and initial guesses. PROXNLP was unable to solve the Push T problem within a tractable amount of time due to its larger problem size, so its data is not applicable in the figure.
@@
Quantitative results. For the first four examples, we recorded and summarized the numerical performance of the benchmark solvers under different initial states and initial guesses, as presented in the Table II and Fig. 5. We set a constraint violation threshold of $1\times 10^{-6}$ for each solver when running the experiments, and all final results with feasibility below $1\times 10^{-5}$ were considered successful, with their average constraint violation recorded. The objective function, as previously highlighted, primarily balances tracking error and control efforts. We analyze the performance of four solvers across three examples, focusing on median tracking error, average iteration count, and computation time. To better illustrate the performance and robustness of each solver under different initial guesses, we boxplot the distribution of constraint violations, objective values, and tracking errors in Fig. 5. These plots provide a comprehensive overview of solver behavior across all experiments, including both feasible and infeasible solutions, offering a complete picture of their performance characteristics.
@@
Real-World Push T. To further validate the effectiveness of CRISP in delivering high-quality trajectories, we applied CRISP to a real robotic arm system for the push T task. In our validation experiments, we estimated the position of the T block during each perception cycle, using this as the new initial state for CRISP to solve them online. The solution from CRISP was then applied in an MPC feedback loop. In the physical experiments, we discretized the problem into 10 time steps with intervals of 0.05 s. Under this setup, the average solution time for CRISP was 80 ms, demonstrating robust and efficient performance in pushing the T block to the target position. A visualization of the process for different initial conditions is provided in Fig. 8.
@@
Figure 8: Key frames visualization of the real-world Push T experiments from various initial configurations (each row). Videos of the results can be found on the project website.
@@ B-C Transport
B-D Push T
Contact-Implicit Formulation: Denote the contact point in body frame (located at the COM of T block) as $(c_{x},c_{y})$, we have 8 different contact mode that exhibit their specific complementarity constraints.
$-2l\leq$ $c_{x}\leq 2l,-d_{c}l\leq c_{y}\leq(4-d_{c})l,$ (76)
$\lambda_{1}\neq 0$ $\Rightarrow c_{y}-(4-d_{c})l=0,$ (77)
$\lambda_{2}\neq 0$ $\Rightarrow|c_{x}-2l|+|c_{y}-(4-d_{c})l|+|c_{y}-(3-d_{c})l|-l=0,$ (78)
$\lambda_{3}\neq 0$ $\Rightarrow|c_{x}-2l|+|c_{x}-0.5l|+|c_{y}-(3-d_{c})l|-1.5l=0,$ (79)
$\lambda_{4}\neq 0$ $\Rightarrow|c_{x}-0.5l|+|c_{y}-(3-d_{c})l|+|c_{y}+d_{c}l|-3l=0,$ (80)
$\lambda_{5}\neq 0$ $\Rightarrow|c_{y}+d_{c}l|+|c_{x}+0.5l|+|c_{x}-0.5l|-l=0,$ (81)
$\lambda_{6}\neq 0$ $\Rightarrow|c_{x}+0.5l|+|c_{y}-(3-d_{c})l|+|c_{y}+d_{c}l|-3l=0,$ (82)
$\lambda_{7}\neq 0$ $\Rightarrow|c_{x}+2l|+|c_{x}+0.5l|+|c_{y}-(3-d_{c})l|-1.5l=0,$ (83)
$\lambda_{8}\neq 0$ $\Rightarrow|c_{x}+2l|+|c_{y}-(4-d_{c})l|+|c_{y}-(3-d_{c})l|-l=0.$ (84)
Designate the positive direction of contact forces as rightward for the $x$ axis and upward for the $y$ axis. And introduce the following slack variables $v$ and $w$ complementing each other to smooth the absolute operation ||$\cdot$|| ⋅ | as we did in the transport example:
$v_{1}-w_{1}$ $=c_{x}-2l,$ (85)
$v_{2}-w_{2}$ $=c_{y}-(4-d_{c})l,$ (86)
$v_{3}-w_{3}$ $=c_{y}-(3-d_{c})l,$ (87)
$v_{4}-w_{4}$ $=c_{x}-0.5l,$ (88)
$v_{5}-w_{5}$ $=c_{y}+d_{c}l,$ (89)
$v_{6}-w_{6}$ $=c_{x}+0.5l,$ (90)
$v_{7}-w_{7}$ $=c_{x}+2l,$ (91)
$0\leq v_{1}$ $\perp w_{1}\geq 0,$ (92)
$0\leq v_{2}$ $\perp w_{2}\geq 0,$ (93)
$0\leq v_{3}$ $\perp w_{3}\geq 0,$ (94)
$0\leq v_{4}$ $\perp w_{4}\geq 0,$ (95)
$0\leq v_{5}$ $\perp w_{5}\geq 0,$ (96)
$0\leq v_{6}$ $\perp w_{6}\geq 0,$ (97)
$0\leq v_{7}$ $\perp w_{7}\geq 0,$ (98)
$0\leq v_{8}$ $\perp w_{8}\geq 0,$ (99)
which can be equivalently written as:
$v_{1}+w_{1}$ $=|c_{x}-2l|,$ (100)
$v_{2}+w_{2}$ $=|c_{y}-(4-d_{c})l|,$ (101)
$v_{3}+w_{3}$ $=|c_{y}-(3-d_{c})l|,$ (102)
$v_{4}+w_{4}$ $=|c_{x}-0.5l|,$ (103)
$v_{5}+w_{5}$ $=|c_{y}+d_{c}l|,$ (104)
$v_{6}+w_{6}$ $=|c_{x}+0.5l|,$ (105)
$v_{7}+w_{7}$ $=|c_{x}+2l|,$ (106)
Then we have the complementarity constraints:
00 0 $\leq-\lambda_{1}\perp(4-d_{c})l-c_{y}\geq 0,$ (108)
00 0 $\leq-\lambda_{2}\perp v_{1}+w_{1}+v_{2}+w_{2}+v_{3}+w_{3}-l\geq 0,$ (109)
00 0 $\leq\lambda_{3}\perp v_{1}+w_{1}+v_{3}+w_{3}+v_{4}+w_{4}-1.5l\geq 0,$ (110)
00 0 $\leq-\lambda_{4}\perp v_{3}+w_{3}+v_{4}+w_{4}+v_{5}+w_{5}-3l\geq 0,$ (111)
00 0 $\leq\lambda_{5}\perp v_{4}+w_{4}+v_{5}+w_{5}+v_{6}+w_{6}-l\geq 0,$ (112)
00 0 $\leq\lambda_{6}\perp v_{3}+w_{3}+v_{5}+w_{5}+v_{6}+w_{6}-3l\geq 0,$ (113)
00 0 $\leq\lambda_{7}\perp v_{3}+w_{3}+v_{6}+w_{6}+v_{7}+w_{7}-1.5l\geq 0,$ (114)
00 0 $\leq\lambda_{8}\perp v_{2}+w_{2}+v_{3}+w_{3}+v_{7}+w_{7}-l\geq 0.$ (115)
Furthermore, if we want to avoid simultaneous application of forces on adjacent edges at corners, we introduce additional complementarity constraints. These constraints ensure that at any given time, only one force can be active:
$\lambda_{i}\perp\lambda_{j},\forall i,j=1\ldots 8,\text{and}\,i\neq j.$ (116)
With the complementarity constraints derived above, we can write the unified dynamics regardless of the contact point.
$\dot{p}_{x}$ $=\frac{1}{\mu mg}.\left[(\lambda_{2}+\lambda_{4}+\lambda_{6}+\lambda_{8})\cos\theta-(\lambda_{1}+\lambda_{3}+\lambda_{5}+\lambda_{7})\sin\theta\right],$ (117)
$\dot{p}_{y}$ $=\frac{1}{\mu mg}.\left[(\lambda_{2}+\lambda_{4}+\lambda_{6}+\lambda_{8})\sin\theta+(\lambda_{1}+\lambda_{3}+\lambda_{5}+\lambda_{7})\cos\theta\right],$ (118)
$\dot{\theta}$ $=\frac{1}{cr\mu mg}.\left[-c_{y}(\lambda_{2}+\lambda_{4}+\lambda_{6}+\lambda_{8})+c_{x}(\lambda_{1}+\lambda_{3}+\lambda_{5}+\lambda_{7})\right],$ (119)
where $c\in[0,1]$ is the integration constant of the box, and $r$ is the characteristic distance, typically chosen as the max contact distance.
B-E Hopper
@@ Appendix C Visualization of Numerical Results
Figure 9: Visualization of more trajectories solved by CRISP in cartpole with soft walls (row 1-2), push box (row 3-4), payload transport (row 5-6), push T (row 7-8) under different initial conditions.